OpenStudy (anonymous):

please help me i am so lost!?! What is the solution of the equation? (equation below in comment)

OpenStudy (anonymous):

\[3\sqrt[5]{(x+2)}^3 +3=27\]

OpenStudy (anonymous):

3(x + 2)^(3/5) + 3 = 27 -> 3(x + 2)^(3/5) = 24 (x + 2)^(3/5) = 8 -> (x + 2)^3 = 8^5 = 2^15 x + 2 = 2^5 = 32 -> x = 30

OpenStudy (anonymous):

The first step, we just subtract 3 from each side, then divide each side by 3, then raise each side to the 5th power, then take the cubic root, then subtract 2.